# Diagonalization $of$ Polynomial-Time Deterministic Turing Machines via Nondeterministic Turing Machines

### v31

@@ 5.1 The $2$nd Simpler Proof
then there must exists at least a fixed $i\in\mathbb{N}_{1}$ such that
@@
steps (or, within time $O(n^{i})$), which clear is a contradiction. We thus can claim that such an $i$ can not be found. Equivalently,
@@
for all $i\in\mathbb{N}_{1}$, which, together with the relations (1) and (2), further imply
@@
Remark 5.1
Comparing the above two proofs of $L_{d}\in\mathcal{NP}$, we are in favour of the second one because it is more concise and simpler than the first proof.
In addition, when an earlier version of this manuscript previously submitted to a journal, we got the feedback that our simulating machine $M_{0}$ does not run in polynomial time whose subtext asserts that
@@
(the submitted version of proof is in Remark 4 of v9 or more clear version is the Lemma 3 in v10). With regard to those comments, we present Theorem 21 most clearly and most rigorously to show that
@@
So the skeptic can carefully examine the proof of Theorem 21 and then point out whether the proof is correct or not, and further she/he can point out why our proof of Theorem 21 is wrong.242424 In fact, the skeptic should give counter-examples to the proof of Theorem 21 if she/he wants to question Theorem 21. For example, to prove that all nondeterministic Turing machines accepting the language $L_{d}$ require $\Omega(c^{n})$ steps, for some constant $c>1$.
@@
mathematically, but some readers are with such a question: Can we find a fixed constant
@@
such that the nondeterministic Turing machine $M_{0}$ runs within polynomial-time
@@
$\text{order}(T_{i})$
be the order of $T_{i}$, i.e., $\text{order}(T_{i})$ is the degree of minimal polynomial of $T_{i}$. For example, if $T_{i}$ is a deterministic $n^{l}+l$ time-bounded Turing machine, then
$\text{order}(T_{i})=l.$
@@
$m=\max\{\text{order}(T_{1}),\text{order}(T_{2}),\cdots\},$
@@
is the minimal polynomial of $M_{0}$. But, here comes the key question, i.e., can we find such a fixed constant $z$ in $\mathbb{N}_{1}$ such that
@@
5.2 Proof of Theorem 1
Now we are at the point to give the proof of Theorem 1:
Proof. It is obvious that Theorem 1 is an immediate consequence of Theorem 9 and Theorem 21.
Remark 5.2
Originally, we call $M_{0}$ a polynomial-time nondeterministic Turing machine. But, complexity theorist Lance Fortnow [For21]252525 We received criticisms [For21] from Lance Fortnow after posting the manuscript of this work on the arXiv. argues that $M_{0}$ does not run in polynomial time because it runs in time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$. He thinks the notion of a polynomial-time machine is a fixed mathematical definition, i.e., call a machine runs in polynomial-time iff the machine runs in time $O(n^{c})$ for some fixed constant $c>0$. For this, see special thanks expressed in acknowledgements subsection 8.1.
However, on the other hand, as we have shown in Theorem 21, the language $L_{d}$ accepted by $M_{0}$ is indeed in $\mathcal{NP}$, which is sufficient for our discussions. In a nutshell, what we are interested in is that to construct a language
@@ 5.2 Proof of Theorem 1
Hence, the distinction between $M_{0}$ running within time $O(n^{c})$ for some fixed $c$ and $M_{0}$ running within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$ is unimportant in our setting. In fact, there is no machine can run within time $O(n^{c})$ for some fixed $c>0$ accepting the $L_{d}$ because the mathematicians acknowledge that $\mathbb{N}_{1}$ is not bounded from above.262626 In fact, we can imagine if $\mathbb{N}_{1}$ is bounded from above, i.e., $i\leq b\text{ for each }i\in\mathbb{N}_{1},$ then $M_{0}$ is a nondeterministic $n^{b}+b$ time-bounded Turing machine (i.e., it runs within time $O(n^{b})$) where $b$ is the “upper-bound” of $\mathbb{N}_{1}$, from which we have the intuition that $L_{d}\in\mathcal{NP}$.
6 A Language $L_{s}\in\mathcal{P}$ That’s Similar to But Different from $L_{d}$
If Theorem 21 were not proved or not discovered, some experts do not regard that $L_{d}\in\mathcal{NP}$. For example, see the opening descriptions of Section 5. But, we should emphasize that, even though $M_{0}$ does not run within time $O(n^{c})$ for fixed $c>0$, $L_{d}$ is indeed in $\mathcal{NP}$. Let us take a further example which resembles that $L_{d}$ is in $\mathcal{NP}$, i.e., we are going to construct a language $L_{s}\in\mathcal{P}$ with the properties that it also needs a specific deterministic Turing machine running within time $O(n^{m})$ for all $m\in\mathbb{N}_{1}$ to accept. Our next result looks like:
6.1 Simulation of Polynomial-Time DTMs
Theorem 12
@@ 6.1 Simulation of Polynomial-Time DTMs
Proof. Let $M^{\prime}_{0}$ be a four-tape deterministic Turing machine which operates as follows on an input string $x$ of length of $n$.
  1. 1.
    $M^{\prime}_{0}$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some single-tape polynomial-time deterministic Turing machine $\widehat{M}_{j}$ for some $j$ then rejects, else determines $t$, the number of tape symbols used by $\widehat{M}_{j}$; $s$, its number of states; and $k$, its order. The third tape of $M^{\prime}_{0}$ can be used as “scratch” memory to calculate $t$.
  2. 2.
@@
    cells each, the blocks being separated by single cell holding a marker ##\##, i.e., there are
@@
  4. 4.
    $M^{\prime}_{0}$ simulates $\widehat{M}_{j}$, using tape $1$, its input tape, to determine the moves of $\widehat{M}_{j}$ and using tape $2$ to simulate the tape of $\widehat{M}_{j}$. The moves of $\widehat{M}_{j}$ are counted in binary in the block of tape $3$, and tape $4$ is used to hold the state of $\widehat{M}_{j}$. If the counter on tape $3$ overflows, $M^{\prime}_{0}$ turns itself off mandatorily and rejects, else $M^{\prime}_{0}$ accepts if and only if $\widehat{M}_{j}$ accepts.
By the arguments similar to those in proof of Theorem 10, it is not hard to show that for any input $w$ to $M^{\prime}_{0}$, there is a corresponding positive integer $i_{w}\in\mathbb{N}_{1}$ such that $M^{\prime}_{0}$ runs at most $|w|^{i_{w}+1}$ steps, thus demonstrating that the deterministic Turing machine $M^{\prime}_{0}$ described above is of time complexity $O(n^{k})$ for any $k\in\mathbb{N}_{1}$ and, of course, accepts some language $L_{s}$. This completes the proof.
6.2 Proof of $L_{s}$ in $\mathcal{P}$
@@
Lemma 6.1
$L_{s}\in\mathcal{P}$ where $L_{s}$ accepted by $M_{0}^{\prime}$ within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$.
Proof. By Corollary 11, the deterministic Turing machine $M^{\prime}_{0}$ is equivalent to a single-tape deterministic Turing machine being of time complexity $O(n^{2k})$ for any $k\in\mathbb{N}_{1}$.
@@
$L_{s}^{i}\subseteq L_{s}^{i+1},\quad\text{for each fixed $i\in\mathbb{N}_{1}$}$
@@
Note further that for any fixed $i\in\mathbb{N}_{1}$, $L_{s}^{i}$ is the language accepted by $M^{\prime}_{0}$ within time $O(n^{i})$. In other words, $L_{s}^{i}$ is accepted by a single-tape deterministic Turing machine $\mathcal{M}^{\prime}_{0}$ of time complexity $O(n^{2k})$ for any $k\in\mathbb{N}_{1}$, which is equivalent to $M^{\prime}_{0}$, within $O(n^{2i})$ steps, we thus obtain
@@
From (4), (5) and (6) we deduce that
@@
6.3 Proof of Theorem 4
Although the language $L_{s}$ is in $\mathcal{P}$ as shown above, we conjecture that there is no deterministic Turing machine running within time $O(n^{c})$ for a fixed $c>0$ that can accept it. Of course, by our constructions of machines $M_{0}$ in Theorem 9 and $M_{0}^{\prime}$ in Theorem 12, it is easy to see that $L_{d}\neq L_{s}$. 272727 $L_{d}$ and $L_{s}$ denote diagonalization language and simulation language, respectively. Furthermore, the $\mathcal{M}^{\prime}_{0}$, which is equivalent to $M^{\prime}_{0}$, is also in the enumeration.
Now the proof of Theorem 4 can be made naturally as follows: Proof of Theorem 4. It clearly follows from Theorem 12 and Lemma 6.1. The proof is completed.
Remark 6.1
Can we now deny that the language $L_{s}$ is not in $\mathcal{P}$ due to that the machine $M^{\prime}_{0}$ accepting it runs within $O(n^{k})$ for all $k\in\mathbb{N}_{1}$? Of course, we can not, because the language $L_{s}$ accepted by $M^{\prime}_{0}$ is exactly the union set of all languages in $\mathcal{P}$ but not others. Moreover, still denote the set of all the polynomial-time deterministic Turing machines by $\{(M,k)\}$, then
@@ 6.3 Proof of Theorem 4
7 Is “Relativization Barrier” Really A Barrier
After the preliminary manuscript posted on the online preprint repository, the author received doubts from other an enthusiastic reader to the problem discussed in this work, i.e., any method that leads to $\mathcal{P}\neq\mathcal{NP}$ must overcome the aforementioned “Relativization Barrier”. Although we have forgotten who emailed the author and the original email has flown away, we thank her/him for her/his interest in the author’s manuscript. Here, we systematically discuss this question in response to that enthusiast.
@@ 7 Is “Relativization Barrier” Really A Barrier
which can be repeated arbitrarily often with different queries. If $O$ is a difficult language (say, one that cannot be decided in polynomial time, or is even undecidable), then this oracle gives the Turing machine additional power. We first quote its formal definition as follows:
Definition 7.1 (Cf. [AB09], Deterministic Oracle Turing machines)
A deterministic oracle Turing machine is a deterministic Turing machine $M$ that has a special read-write tape we call $M$’s oracle tape and three special states $q_{query}$, $q_{yes}$, $q_{no}$. To execute $M$, we specify in addition to the input a language $O\subseteq\{0,1\}^{*}$ that is used as the oracle for $M$. Whenever during the execution $M$ enters the state $q_{query}$, the machine moves into the state $q_{yes}$ if $w\in O$ and $q_{no}$ if $w\not\in O$, where $w$ denotes the contents of the special oracle tape. Note that, regardless of the choice of $O$, a membership query to $O$ counts only as single computation step. If $M$ is an oracle machine, $O\subseteq\{0,1\}^{*}$ a language, and $x\in\{0,1\}^{*}$, then we denote the output of $M$ on input $x$ and with oracle $O$ by $M^{O}(x)$.
@@
$\text{NTIME}^{X}[T(n)].$
The notation $\mathcal{P}^{X}$ and $\mathcal{NP}^{X}$ are defined to be the class of languages:
$\mathcal{P}^{X}=\bigcup_{k\in\mathbb{N}_{1}}\text{DTIME}^{X}[n^{k}]$
@@
$\mathcal{NP}^{X}=\bigcup_{k\in\mathbb{N}_{1}}\text{NTIME}^{X}[n^{k}].$
7.1 Proof of Theorem 6
In this subsection, we prove our main result of the Theorem 6. Before starting, we should first remind the reader that the Theorem 15 is an important prerequisite for proving Theorem 1:
We assume that (I) the polynomial-time deterministic (nondeterministic) oracle Turing machines can be effectively represented as strings; and further suppose that (II) there are universal nondeterministic oracle Turing machine that can simulate any other and flip the answer of other deterministic oracle Turing machines; and lastly suppose that (III) the simulation can be done within in time

### v33

@@ 5.1. The $2$nd Simpler Proof
then there must exist at least a fixed $i\in\mathbb{N}_{1}$ such that
@@
steps (or, within time $O(n^{i})$), which clearly is a contradiction. We thus can claim that such an $i$ can not be found. Equivalently,
@@
for all $i\in\mathbb{N}_{1}$, which, together with the relations (1) and (LABEL:eq2), further implies
@@
Remark 5.3.
Comparing the above two proofs of $L_{d}\in\mathcal{NP}$, we are in favor of the second one because it is more concise and simpler than the first proof.
In addition, when an earlier version of this manuscript was previously submitted to a journal, we got the feedback that our simulating machine $M_{0}$ does not run in polynomial time, i.e., it asserts that
@@
. With regard to those comments, we present Theorem 21 most clearly and most rigorously to show that
@@
So one can carefully examine the proof of Theorem 21 and then point out whether the proof is correct or not, and further, she/he can point out why our proof of Theorem 21 is wrong.242424 One can try to present counter-examples to the proof of Theorem 21 or to prove that all nondeterministic Turing machines accepting the language $L_{d}$ require $\Omega(c^{n})$ steps, for some constant $c>1$.
@@
mathematically, some readers are left with such a question: Can we find a fixed constant
@@
such that the nondeterministic Turing machine $M_{0}$ runs within polynomial time
@@
${\rm order}(T_{i})$
be the order of $T_{i}$, i.e., ${\rm order}(T_{i})$ is the degree of the minimal polynomial of $T_{i}$. For example, if $T_{i}$ is a deterministic $n^{l}+l$ time-bounded Turing machine, then
${\rm order}(T_{i})=l.$
@@
$m=\max\{{\rm order}(T_{1}),{\rm order}(T_{2}),\cdots\},$
@@
is the minimal polynomial of $M_{0}$. But here comes the key question, i.e., can we find such a fixed constant $z$ in $\mathbb{N}_{1}$ such that
@@
5.2. Proof of Theorem 1.1
Now we are at the point to give the proof of Theorem 1.1:
Proof. It is obvious that Theorem 1.1 is an immediate consequence of Theorem 4.1 and Theorem 21.
Remark 5.4.
Originally, we call $M_{0}$ a polynomial-time nondeterministic Turing machine. But complexity theorist Lance Fortnow [For21]252525 We received valuable criticisms [For21] from Lance Fortnow after posting the manuscript of this work on the arXiv. argues that $M_{0}$ does not run in polynomial time because it runs in time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$. He thinks the notion of a polynomial-time machine is a fixed mathematical definition, i.e., call a machine running in polynomial time iff the machine runs in time $O(n^{c})$ for some fixed constant $c>0$. For this, see special thanks expressed in acknowledgements subsection LABEL:sec:acks.
However, on the other hand, as we have shown in Theorem 21, the language $L_{d}$ accepted by $M_{0}$ is indeed in $\mathcal{NP}$, which is sufficient for our discussions. In a nutshell, what we are interested in is constructing a language
@@
Hence, the distinction between $M_{0}$ running within time $O(n^{c})$ for some fixed $c$ and $M_{0}$ running within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$ is unimportant in our setting. In fact, there is no machine that can run within time $O(n^{c})$ for some fixed $c>0$ accepting the $L_{d}$ because the mathematicians acknowledge that $\mathbb{N}_{1}$ is not bounded from above.262626 In fact, we can imagine if $\mathbb{N}_{1}$ is bounded from above, i.e., $i\leq b\text{ for each }i\in\mathbb{N}_{1},$ then $M_{0}$ is a nondeterministic $n^{b}+b$ time-bounded Turing machine (i.e., it runs within time $O(n^{b})$) where $b$ is the “upper-bound” of $\mathbb{N}_{1}$, from which we have the intuition that $L_{d}\in\mathcal{NP}$.
If Theorem 21 were not proved or not discovered, some experts would not regard that $L_{d}\in\mathcal{NP}$. For example, see the opening descriptions of Section 5. But, we should emphasize that, even though $M_{0}$ does not run within time $O(n^{c})$ for fixed $c>0$, $L_{d}$ is indeed in $\mathcal{NP}$. Let us take a further example that resembles that $L_{d}$ is in $\mathcal{NP}$, i.e., we are going to construct a language $L_{s}\in\mathcal{P}$ with the properties that it also needs a specific deterministic Turing machine running within time $O(n^{m})$ for all $m\in\mathbb{N}_{1}$ to accept. Our next result looks like the following.
6.1. Simulation of Polynomial-Time DTMs
Theorem 6.1.
@@ 6.1. Simulation of Polynomial-Time DTMs
Proof 6.2.
Let $M^{\prime}_{0}$ be a four-tape deterministic Turing machine that operates as follows on an input string $x$ of length $n$.
  1. (1)
    $M^{\prime}_{0}$ decodes the tuple encoded by $x$. If $x$ is not the encoding of some single-tape polynomial-time deterministic Turing machine $\widehat{M}_{j}$ for some $j$ then rejects; else determine $t$, the number of tape symbols used by $\widehat{M}_{j}$; $s$, its number of states; and $k$, its order. The third tape of $M^{\prime}_{0}$ can be used as “scratch” memory to calculate $t$.
  2. (2)
@@
    cells each, the blocks being separated by a single cell holding a marker ##\##, i.e., there are
@@
  4. (4)
    $M^{\prime}_{0}$ simulates $\widehat{M}_{j}$, using tape $1$, its input tape, to determine the moves of $\widehat{M}_{j}$ and using tape $2$ to simulate the tape of $\widehat{M}_{j}$. The moves of $\widehat{M}_{j}$ are counted in binary in the block of tape $3$, and tape $4$ is used to hold the state of $\widehat{M}_{j}$. If the counter on tape $3$ overflows, $M^{\prime}_{0}$ turns itself off mandatorily and rejects; else $M^{\prime}_{0}$ accepts if and only if $\widehat{M}_{j}$ accepts.
By arguments similar to those in the proof of Theorem 4.4, it is not hard to show that for any input $w$ to $M^{\prime}_{0}$, there is a corresponding positive integer $i_{w}\in\mathbb{N}_{1}$ such that $M^{\prime}_{0}$ runs at most $|w|^{i_{w}+1}$ steps, thus demonstrating that the deterministic Turing machine $M^{\prime}_{0}$ described above is of time complexity $O(n^{k})$ for any $k\in\mathbb{N}_{1}$ and, of course, accepts some language $L_{s}$. This completes the proof.
6.2. Proof of $L_{s}$ in $\mathcal{P}$
@@
Lemma 6.3.
$L_{s}\in\mathcal{P}$ where $L_{s}$ is accepted by $M_{0}^{\prime}$ within time $O(n^{k})$ for any $k\in\mathbb{N}_{1}$.
Proof 6.4.
By Corollary 11, the deterministic Turing machine $M^{\prime}_{0}$ is equivalent to a single-tape deterministic Turing machine being of time complexity $O(n^{2k})$ for any $k\in\mathbb{N}_{1}$.
@@
Similarly, the above definition of language $L_{s}^{i}$ technically can be done by adding a new tape to $M_{0}^{\prime}$ as a counter to count up to
$n^{i+1}$
for a fixed $i\in\mathbb{N}_{1}$, meaning that $M_{0}^{\prime}$ turns itself off when the counter of tape $3$ exceeds
$n^{k+1}$
or the counter of the newly added tape exceeds
$n^{i+1}.$
@@
$L_{s}^{i}\subseteq L_{s}^{i+1},\quad\text{for each fixed $i\in\mathbb{N}_{1}$.}$
@@
Note further that for any fixed $i\in\mathbb{N}_{1}$, $L_{s}^{i}$ is the language accepted by $M^{\prime}_{0}$ running within time $O(n^{i})$. In other words, $L_{s}^{i}$ is accepted by a single-tape deterministic Turing machine $N_{0}$ of time complexity $O(n^{2i})$, which is equivalent to $M^{\prime}_{0}$ running within time $O(n^{i})$. We thus obtain that
@@
From (4), (5), and (6), we deduce that
@@
6.3. Proof of Theorem 1.4
Although the language $L_{s}$ is in $\mathcal{P}$ as shown above, we conjecture that there is no deterministic Turing machine running within time $O(n^{c})$ for a fixed $c>0$ that can accept it. Of course, by our constructions of machines $M_{0}$ in Theorem 4.1 and $M_{0}^{\prime}$ in Theorem 6.1, it is easy to see that $L_{d}\neq L_{s}$. 272727 $L_{d}$ and $L_{s}$ denote diagonalization language and simulation language, respectively.
Now the proof of Theorem 1.4 can be made naturally as follows: Proof of Theorem 1.4. It clearly follows from Theorem 6.1 and Lemma 6.3. The proof is completed.
Remark 6.5.
Can we now deny that the language $L_{s}$ is not in $\mathcal{P}$ due to the machine $M^{\prime}_{0}$ accepting it running within $O(n^{k})$ for all $k\in\mathbb{N}_{1}$? Of course, we can not, because the language $L_{s}$ accepted by $M^{\prime}_{0}$ is in $\mathcal{P}$ but not others.282828 In fact, we can construct a universal deterministic Turing machine running within time $O(n^{i})$ for any $i\in\mathbb{N}_{1}$ accepting the language $L^{\prime}_{s}=\{(\langle M_{i}\rangle,w)\,|\,w\in L(M_{i})\},$ where $\langle M_{i}\rangle$ is the binary string representation of some polynomial-time deterministic Turing machine taken from the set $\{(M,k)\}$, and $L(M_{i})$ denotes the language accepted by $M_{i}$, and the tuple of $(\langle M_{i}\rangle,w)$ can be seen as a pairing function, which is similar to [DK14]; see Proposition 1.17, p. 26 in [DK14]. Moreover, still denote the set of all the polynomial-time deterministic Turing machines by $\{(M,k)\}$, then
@@
7. Breaking the “Relativization Barrier”
After the preliminary manuscript was posted on the online preprint repository, the author received doubts from another enthusiastic reader about the problem discussed in this work, i.e., any method that leads to $\mathcal{P}\neq\mathcal{NP}$ must overcome the aforementioned “Relativization Barrier.” Although we have forgotten who emailed the author and the original email has flown away, we thank her/him for her/his interest in the author’s manuscript. Here, we systematically discuss this question in response to that enthusiast.
@@ 7. Breaking the “Relativization Barrier”
which can be repeated arbitrarily often with different queries. If $O$ is a difficult language (say, one that cannot be decided in polynomial time or is even undecidable), then this oracle gives the Turing machine additional power. We first quote its formal definition as follows:
Definition 7.1 (Cf. [AB09], Deterministic Oracle Turing machines).
A deterministic oracle Turing machine is a deterministic Turing machine $M$ that has a special read-write tape we call $M$’s oracle tape and three special states $q_{query}$, $q_{yes}$, and $q_{no}$. To execute $M$, we specify in addition to the input a language $O\subseteq\{0,1\}^{*}$ that is used as the oracle for $M$. Whenever during the execution $M$ enters the state $q_{query}$, the machine moves into the state $q_{yes}$ if $w\in O$ and $q_{no}$ if $w\not\in O$, where $w$ denotes the contents of the special oracle tape. Note that, regardless of the choice of $O$, a membership query to $O$ counts only as a single computation step. If $M$ is an oracle machine, $O\subseteq\{0,1\}^{*}$ a language, and $x\in\{0,1\}^{*}$, then we denote the output of $M$ on input $x$ and with oracle $O$ by $M^{O}(x)$.
@@
${\rm NTIME}^{X}[T(n)].$
The notation $\mathcal{P}^{X}$ and $\mathcal{NP}^{X}$ is defined respectively to be the class of languages:
$\mathcal{P}^{X}=\bigcup_{k\in\mathbb{N}_{1}}{\rm DTIME}^{X}[n^{k}]$
@@
$\mathcal{NP}^{X}=\bigcup_{k\in\mathbb{N}_{1}}{\rm NTIME}^{X}[n^{k}].$
7.1. Proof of Theorem 6
In this subsection, we prove our main result of Theorem 6. Before starting, we should first remind the reader that Theorem 15 is an important prerequisite for proving Theorem 1.1:
We assume that (I) the polynomial-time deterministic (nondeterministic) oracle Turing machines can be effectively represented as strings; and further suppose that (II) there are universal nondeterministic oracle Turing machines that can simulate any other and flip the answer of other deterministic oracle Turing machines; and lastly suppose that (III) the simulation can be done within time